# On long-duration storage, weather uncertainty and limited foresight

### v1

@@ Acknowledgments
Based on the forward pass, we obtain a vector of outgoing states $\{x_{t}^{k}\}$. In the backward pass we start in stage $T-1$ since the stage $T$ does not have a cost-to-go term. In stage $T-1$, we take candidate state ${x_{T-1}^{k}}$ and solve the subproblem in stage $T$, ${Q}_{T}(x^{k}_{T-1},\xi_{T,i})$ for all $\xi_{T,i}\in\Xi_{T}$. For each weather realization $i$, we compute the dual value $\rho^{k}_{T}(x_{T-1}^{k},\xi_{T,i})$ and evaluate ${Q}_{T}(x^{k}_{T-1},\xi_{T,i})$ to construct a new cut in $T-1$ according to Eq. SI.3e.
Next we move to stage $T-2$, take the candidate outgoing state $x_{T-2}^{k}$ and solve the approximate subproblem for $T-1$, $\hat{Q}_{T-1}(x^{k}_{T-2},\xi_{T-1,i})$ for all $\xi_{T-1,i}\in\Xi_{T-1}$. Again, we compute dual variables $\rho_{T-1}^{k}(x_{T-2}^{k},\xi_{T-1,i})$ and cost-to-go $\hat{Q}_{T-1}(x^{k}_{T-2},\xi_{T-1,i})$ for all weather realizations and a add a new cut to the subproblem in stage $T-2$. Note that in this stage the next stage’s subproblem is the approximate model $\hat{Q}^{k}_{T-1}$ that has already been updated with cuts in the previous step of the backward pass. We continue the procedure until stage 00.
@@
Stochastic dual dynamic programming does not have a deterministic upper bound as we cannot evaluate the entire state space (recall that the number of scenarios is a very large number $N$). We can simulate a Monte Carlo estimator of the upper bound instead. To this end, we simulate a collection so sample paths ${\xi_{i}}\ i\in\mathcal{M}$ where $|\mathcal{M}|<<N$ and simulate forward passes on these paths to calculate a Monte Carlo estimator for the upper bound,
@@
The classical termination rule for SDDP would be to stop once the lower bound is within the confidence interval of the upper bound [pereira_multi-stage_1991]. However, this rule has its weaknesses [dowson_policy_2020] and termination is an open research topic.
@@
To facilitate a comparison of the dispatch behavior between the Limited and the Perfect Foresight models, we use the resulting policy to simulate Limited Foresight in the historical weather years in the sample. First, note that the set of historical weather years $\Xi^{hist}$ is a comparably small subset of $\Xi$ with $N^{hist}=35$ in the case study above. For each historical weather year $i\in\Xi^{hist}$, we compute the solution to a forward pass. The capacities remain the same for all historical weather years and the dispatch decisions results from the trained policy $\{y^{K}_{t}(\xi_{t,i})\}_{t=1}^{T}$.
@@
Lastly, we discuss how to we derive the marginal storage values or long-duration energy storage bidding curves from the trained model in in Section 4.3.
@@
Figure SI.1: Objective value convergence: Convergence of the lower bound corresponding to the model’s first stage objective value and the 2000 iterations rolling mean simulations as an estimator for the upper bound for all 9 model specifications.
@@
Figure SI.6: Optimal capacities with constrained outside option
Figure SI.7: Optimal capacities with unlimited outside option
@@
Figure SI.9: Mean hourly electrolysis capacity factor by time of day in No Imports
@@
Table SI.2: Correlation of pv and wind generation with PEM electrolysis by scenario
Figure SI.10: Daily electrolysis and reconversion capacity factors across all weather years
@@
Table SI.6: Aggregate energy demands
SI.10 Stagewise independence test
In this section we describe the approach to compute the autocorrelation functions to justify the stagewise independence and rationalize the foresight assumption. We take four-hourly data for each of the weather variables of interest. For a given weather variable $x$ we compute the monthly average value given by,
@@
where $\sigma$ is the standard deviation of the monthly time series. Figure 2 and Figure SI.15 plot these autocorrelations as a function of $k$.
Figure SI.15: Autocorrelation functions for weekly stage length

### v2

@@ Acknowledgments
Based on the forward pass, we obtain a vector of outgoing states $\{x_{t}^{k}\}$. In the backward pass, we start in stage $T-1$ since the stage $T$ does not have a cost-to-go term. In stage $T-1$, we take candidate state ${x_{T-1}^{k}}$ and solve the subproblem in stage $T$, ${Q}_{T}(x^{k}_{T-1},\xi_{T,i})$ for all $\xi_{T,i}\in\Xi_{T}$. For each weather realization $i$, we compute the dual value $\rho^{k}_{T}(x_{T-1}^{k},\xi_{T,i})$ and evaluate ${Q}_{T}(x^{k}_{T-1},\xi_{T,i})$ to construct a new cut in $T-1$ according to Eq. SI.3e.
Next we move to stage $T-2$, take the candidate outgoing state $x_{T-2}^{k}$ and solve the approximate subproblem for $T-1$, $\hat{Q}_{T-1}(x^{k}_{T-2},\xi_{T-1,i})$ for all $\xi_{T-1,i}\in\Xi_{T-1}$. Again, we compute dual variables $\rho_{T-1}^{k}(x_{T-2}^{k},\xi_{T-1,i})$ and cost-to-go $\hat{Q}_{T-1}(x^{k}_{T-2},\xi_{T-1,i})$ for all weather realizations and a add a new cut to the subproblem in stage $T-2$. Note that in this stage, the next stage’s subproblem is the approximate model $\hat{Q}^{k}_{T-1}$ that has already been updated with cuts in the previous step of the backward pass. We continue the procedure until stage 00.
@@
Stochastic dual dynamic programming does not have a deterministic upper bound as we cannot evaluate the entire state space (recall that the number of scenarios is a very large number $N$). We can simulate a Monte Carlo estimator of the upper bound instead. To this end, we simulate a collection of sample paths ${\xi_{i}},\ i\in\mathcal{M}$, where $|\mathcal{M}|<<N$ and simulate forward passes on these paths to calculate a Monte Carlo estimator for the upper bound,
@@
The classical termination rule for SDDP would be to stop once the lower bound is within the confidence interval of the upper bound [pereira_multi-stage_1991]. However, this rule has its weaknesses [dowson_policy_2020], and termination is an open research topic.
@@
To facilitate a comparison of the dispatch behavior between the Limited and the Perfect Foresight models, we use the resulting policy to simulate Limited Foresight in the historical weather years in the sample. First, note that the set of historical weather years $\Xi^{hist}$ is a comparably small subset of $\Xi$ with $N^{hist}=35$ in the case study above. For each historical weather year $i\in\Xi^{hist}$, we compute the solution to a forward pass. The capacities remain the same for all historical weather years, and the dispatch decisions result from the trained policy $\{y^{K}_{t}(\xi_{t,i})\}_{t=1}^{T}$.
@@
Lastly, we discuss how to derive the marginal storage values or long-duration energy storage bidding curves from the trained model in Section 4.3.
@@
This section provides some details on model convergence for the Limited Foresight model solved with stochastic dual dynamic programming (SDDP). Figure SI.1 shows the convergence of the model’s lower bound and a rolling average of forward pass simulations for all model specifications.
Figure SI.1: Objective value convergence: Convergence of the lower bound corresponding to the model’s first stage objective value and the 100 iterations rolling mean simulations as an estimator for the upper bound for all 9 model specifications.
Table SI.1 provide run times and iteration number for all specifications and Figure SI.2 shows the convergence of the capacity choices in stage 0.
@@
This section presents additional results. Specifically, Figures SI.3 and SI.4 replicate Figure 4.
@@
Figure SI.5 shows the storage trajectories for 35 historical weather years in chronological sequence.
@@
Figures SI.6 and SI.7 show the capacity decisions under Perfect and Limited Foresight for the Constrained Imports and Unlimited Imports scenarios.
Figure SI.6: Optimal capacities - Constrained Imports
Figure SI.7: Optimal capacities - Unlimited Imports
@@
Figure SI.9: Mean hourly electrolysis capacity factor by time of day (No Imports)
@@
Table SI.2: Correlation of pv and wind generation with PEM electrolysis by scenario (No Imports)
Figure SI.10: Daily electrolysis and reconversion capacity factors across all weather years (No Imports)
Figures SI.11 to SI.13 show the realized marginal storage values for the simulation period.
@@
Table SI.6:
Aggregate energy demands
.10 Stagewise independence test
In this section, we describe the approach to computing the autocorrelation functions to justify the stagewise independence and rationalize the foresight assumption. We take four-hourly data for each of the weather variables of interest. For a given weather variable $x$ we compute the monthly average value given by,
@@
where $\sigma$ is the standard deviation of the monthly time series. Figure 2 and Figure 15 plot these autocorrelations as a function of $k$.
Figure 15: Autocorrelation functions for weekly stage length